# Wormhole spacetimes in an expanding universe: energy conditions and future singularities

### v1

@@ I Introduction
An exotic compact object called a wormhole was first proposed in Refs. [1, 2]. Such geometries connect distant regions of the universe or even entirely separate universes [3]. These structures require matter that violates the null energy condition (NEC), and hence all the standard energy conditions, including the weak energy condition (WEC), strong energy condition (SEC) and dominant energy condition (DEC). Such matter is often referred to as exotic matter and typically leads to instabilities. Possible origins of such matter include higher-dimensional effects [4]. Wormholes have attracted renewed attention not only as theoretical curiosities but also as potential probes of fundamental physics.
In this work, we study wormholes embedded in an expanding universe. Although phantom dark energy violates all the energy conditions, the presence of the cosmological fluid responsible for the expansion can modify the total energy density and pressure, thereby relaxing the violation, as discussed in Ref. [5]. A related model was proposed in Ref. [6], where Einstein gravity coupled to two scalar fields generates a wormhole geometry. In this model, the energy conditions are violated due to the energy–momentum tensor of the scalar fields, and one of the fields becomes a ghost. Classically, such a ghost leads to an unbounded kinetic energy, and in quantum theory, it gives rise to negative-norm states, which undermine the consistency of quantum mechanics. However, the ghost can be removed by introducing a Lagrange multiplier field, which imposes a constraint that effectively freezes the scalar dynamics, leading to a non-dynamical but stable configuration.
An extension including a Gauss–Bonnet term was proposed in Ref. [7], leading to time-dependent, or dynamical, wormhole geometries that exist only within a finite time interval. To construct a general spherically symmetric and dynamical geometry, two scalar fields are sufficient. However, to realise an entirely arbitrary geometry, four scalar fields are necessary and sufficient [8, 9]. In this framework, the dynamical wormhole geometries are formulated as a non-linear $\sigma$ model in which the target-space metric is identified with the Ricci tensor of the spacetime, and the four scalar fields correspond to spacetime coordinates. Because of the inherent constraints in this model, the geometry is stable even when the energy conditions are violated.
@@ VII.3 Case 3: w<1w<-1
Here, we input $t_{0}=0$. We plot the above six quantities as functions of $(t/t_{s},\chi/r_{0})$. Fig. 6 shows $\rho$, $\rho+p^{\mathrm{radial}}$, $\rho+p^{\mathrm{angular}}$, $\rho+p^{\mathrm{radial}}+2p^{\mathrm{angular}}$, $\rho-\left|p^{\mathrm{radial}}\right|$ and $\rho-\left|p^{\mathrm{angular}}\right|$ for $r_{0}H_{0}=10^{-2}$ and w=4/3w=-4/3. All the energy conditions are always violated as in the standard phantom universe.
@@
Figure 6: The behaviours of $\rho$ (6(a)), $\rho+p^{\mathrm{radial}}$ (6(b)), $\rho+p^{\mathrm{angular}}$ (6(c)), $\rho+p^{\mathrm{radial}}+2p^{\mathrm{angular}}$ (6(d)), $\rho-\left|p^{\mathrm{radial}}\right|$ (1(e)) and $\rho-\left|p^{\mathrm{angular}}\right|$ (1(f)), when we use the normalisation of $\frac{{H_{0}}^{2}}{\kappa^{2}}$ with $r_{0}H_{0}=10^{-2}$, w=4/3w=-4/3. All the energy conditions are always violated as in the standard phantom universe.
@@ VIII Singular spacetime with wormhole
and $t$ can be identified with the cosmic time in the FLRW universe. With this asymptotic identification, the relation with future singularities becomes manifest. In the following, we demonstrate two particular examples of the spacetime that shows a singularity.
@@ IX Wormhole motivated by Dark Matter Halo
In previous sections, we have studied the wormhole geometry embedded into the FLRW universe and analysed the energy conditions in such a spacetime. However, it is possible to embed a wormhole geometry into another spacetime. There was recently some interest in wormholes corrected by the density profiles of dark matter (DM) halos in galaxies. For instance, Ref. [26] is devoted to wormhole geometry corrected by the Navarro-Frenk-White (NFW) profile [28, 29] and other profiles [30, 31, 32]. In Ref. [26], the wormhole is assumed to alternate the supermassive black hole at the galactic centre. Several other kinds of wormhole spacetime with DM profile accounts have been proposed, see [27] In this section, we briefly review such a wormhole geometry corrected by a DM profile. We demonstrate that such a wormhole may be naturally realised within our four-scalar framework. In other words, the four-scalar framework gives the same effect as DM profile to a wormhole without any account of real DM. In this section, we briefly review a wormhole geometry in [27] as one example and examine it within our four-scalar framework.
@@
We should note that any other kinds of wormhole geometries, based on other DM profiles or anything else, can also be stably realised by using the non-linear $\sigma$ model with four scalar fields given in (1) - (4). We do not need to use any real DM or compensate the real DM by using the four-scalar model.

### v2

@@ I Introduction
An exotic compact object called a wormhole was first proposed in Refs. [1, 2]. Such geometries connect distant regions of the universe or even entirely separate universes [3]. These structures require matter that violates the null energy condition (NEC) and, hence, all the standard energy conditions, including the weak energy condition (WEC), strong energy condition (SEC), and dominant energy condition (DEC). Such matter is often referred to as exotic matter and typically leads to instabilities. Possible origins of such matter include higher-dimensional effects [4]. Wormholes have attracted renewed attention not only as theoretical curiosities but also as potential probes of fundamental physics.
A pioneering work [5] showed the construction of the wormhole solution embedded in the expanding universe satisfying the energy conditions. Moreover, effects of the non-local gravity, the inverse d’Alembert operator in the gravitational sector, can stabilize wormhole solutions [6], and the wormhole spacetime can be supported without exotic matter in the Einstein-Gauss-Bonnet gravity with the four-dimensional limit [7]. Building on numerous earlier results, we aim to investigate the concrete realization of the energy–momentum tensor required to support the wormhole geometry within a self-consistent matter sector. We also demonstrate a novel field-theoretical framework in which the wormhole solution is supported, rather than relying on an effective or phenomenological fluid description.
In this work, we study wormholes embedded in an expanding universe. Although phantom dark energy (DE) violates all the energy conditions, the presence of the cosmological fluid responsible for the expansion can modify the total energy density and pressure, thereby relaxing the violation, as discussed in Ref. [8]. A related model was proposed in Ref. [9], where Einstein gravity coupled to two scalar fields generates a wormhole geometry. In this model, the energy conditions are violated due to the energy–momentum tensor of the scalar fields, and one of the fields becomes a ghost. Classically, such a ghost leads to an unbounded kinetic energy, and in quantum theory, it gives rise to negative-norm states, which undermine the consistency of quantum mechanics. However, the ghost can be removed by introducing a Lagrange multiplier field, which imposes a constraint that effectively freezes the scalar dynamics, leading to a non-dynamical but stable configuration.
An extension including a Gauss–Bonnet term was proposed in Ref. [10], leading to time-dependent, or dynamical, wormhole geometries that exist only within a finite time interval. To construct a general spherically symmetric and dynamical geometry, two scalar fields are sufficient. However, to realise an entirely arbitrary geometry, four scalar fields are necessary and sufficient [11, 12]. In this framework, the dynamical wormhole geometries are formulated as a non-linear $\sigma$ model in which the target-space metric is identified with the Ricci tensor of the spacetime [13], and the four scalar fields correspond to spacetime coordinates. Because of the inherent constraints in this model, the geometry is stable even when the energy conditions are violated. The scalar fields and non-linear $\sigma$ models often appear in the theories originated from higher-dimensional theories like string theory, and the gauge symmetries, including general covariance in higher dimensions, may give constraints as in the non-linear $\sigma$ model.
@@ VII.3 Case 3: w<1w<-1
Here, we input $t_{0}=0$. We plot the above six quantities as functions of $(t/t_{s},\chi/r_{0})$. Fig. 6 shows $\rho$, $\rho+p^{\mathrm{radial}}$, $\rho+p^{\mathrm{angular}}$, $\rho+p^{\mathrm{radial}}+2p^{\mathrm{angular}}$, $\rho-\left|p^{\mathrm{radial}}\right|$ and $\rho-\left|p^{\mathrm{angular}}\right|$ for $r_{0}H_{0}=10^{-2}$ and w=4/3w=-4/3. All the energy conditions are always violated, as in the standard phantom universe.
@@
Figure 6: The behaviours of $\rho$ (6(a)), $\rho+p^{\mathrm{radial}}$ (6(b)), $\rho+p^{\mathrm{angular}}$ (6(c)), $\rho+p^{\mathrm{radial}}+2p^{\mathrm{angular}}$ (6(d)), $\rho-\left|p^{\mathrm{radial}}\right|$ (1(e)) and $\rho-\left|p^{\mathrm{angular}}\right|$ (1(f)), when we use the normalisation of $\frac{{H_{0}}^{2}}{\kappa^{2}}$ with $r_{0}H_{0}=10^{-2}$, w=4/3w=-4/3. All the energy conditions are always violated, as in the standard phantom universe.
VII.4 Recovery of energy conditions
In the previous three subsections, we considered embedding the wormhole into the accelerated expansion of the Universe for three representative cases: quintessence, a cosmological constant, and phantom energy. We find that the contribution of these DE components enters the total energy–momentum tensor through an effective cosmological fluid, whose dominance depends on the relative scales of the wormhole $r_{0}$ and the background expansion $H_{0}$.
For quintessence and the cosmological constant, we find that the energy conditions can be recovered when the DE component becomes dominant in the total energy budget. This corresponds to a regime in which the repulsive effect associated with cosmic acceleration sufficiently contributes to the overall energy and pressure, compensating for the violations that typically arise from the wormhole sector. An important outcome of this analysis is that the recovery of the energy conditions is not universal, but scale-dependent. The energy conditions are evaluated in terms of the ratio of the wormhole radius to the Hubble radius $r_{0}H_{0}$. When the wormhole size is much smaller than the Hubble radius $r_{0}H_{0}\ll 1$, the local gravitational effects associated with the wormhole geometry become relatively stronger, making it more difficult for the cosmic expansion to restore the energy conditions.
In contrast, for the phantom case, the situation is qualitatively different. Since phantom energy intrinsically violates several energy conditions, the inclusion of this component does not lead to an overall recovery of the energy conditions, even when cosmic expansion is taken into account. This distinction highlights that the recovery mechanism is not a generic consequence of cosmic expansion, but rather depends on the physical nature of the cosmological fluid.
@@ VIII Singular spacetime with wormhole
and $t$ can be identified with the cosmic time in the FLRW universe. With this asymptotic identification, the relation with future singularities becomes manifest.
In the following, we demonstrate the embedding of the wormhole into the spacetime, showing a future singularity, in two particular examples. Note that the finite-time future singularity arises from the effective DE behavior induced by the scalar-field sector, rather than from the wormhole geometry itself. That is, the wormhole solution is embedded in a cosmological background driven independently by the dark-energy component arising from the four-scalar fields. We illustrate cosmological scenarios where both phenomena can naturally emerge.
@@ IX Wormhole motivated by Dark Matter Halo
In previous sections, we have studied the wormhole geometry embedded into the FLRW universe and analysed the energy conditions in such a spacetime. However, it is possible to embed a wormhole geometry into another spacetime. There was recently some interest in wormholes corrected by the density profiles of dark matter (DM) halos in galaxies. For instance, Ref. [17] is devoted to wormhole geometry corrected by the Navarro-Frenk-White (NFW) profile [15, 16] and other profiles [19, 18, 29]. In Ref. [17], the wormhole is assumed to alternate the supermassive black hole at the galactic centre. Several other kinds of wormhole spacetime with DM profile accounts have been proposed (see [30]). In this section, we briefly review such a wormhole geometry corrected by a DM profile. We demonstrate that such a wormhole may be naturally realised within our four-scalar framework. In other words, the four-scalar framework gives the same effect as DM profile to a wormhole without any account of real DM. In this section, we briefly review a wormhole geometry in [30] as one example and examine it within our four-scalar framework.
@@
We should note that any other kinds of wormhole geometries, based on other DM profiles or anything else, can also be stably realised by using the four-scalar non-linear $\sigma$ model given in (1) - (4). We do not need to use any real DM or compensate the real DM by using the four-scalar model. Our analysis merely highlights a limitation of the proposed DM halo scenario [17] within the four-scalar non-linear $\sigma$ model, and the absence of a pressureless DM component in this context can be viewed as a constraint on that specific application.